PS2PK(1) PS2PK(1) NAME ps2pk−createsaTeXpkfontfromatype1PostScriptfont SYNOPSIS pk2pk[options]type1[pkfont] options:-v-eenc-Xxres-Eexpansion-Sslant options:-Ppointsize-Rbaseres-raspectratio options:-Yyres-aAFMfile-mmodename DESCRIPTION Thisprogramrendersagiventype1PostScriptfontatagivenpointsize(default10.0points)and resolution(default300dpi)intoaTeXpkfont. Togeneratethepkfontpk2psneedsavalidtype1fontname(forexampleUtopia-Regular,Utopia-Regula.pfborUtopia-Regula.pfa)anditscorrespondingAFM-file(Utopia-Regular.afm).Thepro-gramacceptsboththeMSDOSbinarytype1format(Utopia-Regula.pfb)anditsASCIIequivalent (Utopia-RegularorUtopia-Regula.pfa).Tolocatethetype1fontanditsAFM-fileps2pkwilluse theenvironmentT1INPUTSifitissetotherwiseitsbuiltindefault(the-vflagwillreportwhich filenamesarelookedfor). TheprogramwillmakeapkfontinwhichthecharactercodesarederivedfromtheirAFM(Adobe FontMetrics)definedvalues.Thiscanbeoverruledbyspecifyinganexplicitencodingfileviathe -eoption.TheencodingfileencshouldcontainavalidPostScriptencodingcontainingPostScript nameslike/hyphen.Hereisanincompleteexample(seeafm2tfmforcompletesyntax): %ThisistheECencoding. /ECEncoding[%now256charsfollow %0x00 /grave/acute/circumflex/tilde/dieresis/hungarumlaut /ring/caron/breve/macron/dotaccent/cedilla /ogonek/quotesinglbase/guilsinglleft/guilsinglright (Restoflinesomitted.) %0xF0 /eth/ntilde/ograve/oacute/ocircumflex/otilde /odieresis/oe/oslash/ugrave/uacute/ucircumflex /udieresis/yacute/thorn/germandbls]def Optionsandarguments: -aAFMfileOverrulesthenamethatisusedtolocatetheAFMfile. -eencNameofafilecontaininganencodingscheme(default−theencodingnamedinthe AFM-fileisused).TheactualPostScriptnameoftheencodingschemeiswrittenas a"special"attheendofthepkfont. -EextensionTheextensionfactor(realvalue,default1.0). -mmodename AmodenamesuchaswouldbeusedbyMETAFONT(default"Unknownmode"). -PpointsizeThedesiredpointsize(realvalue,default10.0points). TeX 1 PS2PK(1) PS2PK(1) -RbaseresThedesiredbaseresolution(integervalue,default300dpi).Ifthisdiffersfromthe valueofxres,theappropriatemagnificationwillbeshowninthe"specials"atthe endofthepkfont.Ifpossible,thiswillbeexpressedasamagstep,otherwiseasareal number.Forexample,apkfontwithanxresof329,usingthedefaultbaseresolution of300,willincludethe"special"text"mag=magstep(0.5)"attheendofthefont. -raspectratioThedesiredaspectratio(expression,integerY/integerX,default300/300).Thiswill beshowninthe"specials"attheendofthepkfont.Settingtheaspectratiocreates avalueforyresbutanexplicityresvaluewilltakeprecedence.Argumentmustbe in""ifspacesareleftaround‘/’. -SslantTheslant(realvalue,default0.0). -XxresTheresolutionintheXdirection(integervalue,default300dpi). -YyresTheresolutionintheYdirection(integervalue,defaultthevalueofxres).Ifyres differsfromxres,a"special"textgivingtheaspectratiowillbewrittenattheend ofthepkfont. -vVerboseflag.(Tellswhattheprogramisdoing.) type1ThenameofthePostScripttype1font.ThenameoftheAFM-filewillbecon-structedfromthisnamebyremovingtheextension(ifsuplied)andadding".afm". ThePostScript"FontName"isextractedfromtheAFM-fileandwrittenintoa"spe-cial"attheendofthepkfont. [pkfont]Thenameoftheresultingpkfontcanbeoverruledwiththisname.Thedefault nameofpkfontisderivedfromthebasenameofthetype1font,thepointsizeand xres.Forexample‘ps2pk-P17.28Utopia-Regular’willresultin‘Utopia-Regu-lar17.300pk’.Anexplicitvalueforthenameofthepkfontisnecessarywhenthe type1fontnamealreadyshowsthepointsize,otherwisethepointsizevalueiscatenatedontothepkfontbasenamewhichisusuallynotwhatiswanted. ThefollowingPK"specials"provideaninternalcheckonthecharacteristicsofthepkfont,in accordancewiththerecommendationsoftheTeXWorkingGroupontheTeXDirectoryStandard (TWG-TDS): "fontid=FontName","codingscheme=Encoding","fontfacebyte","mag","mode=(ps2pk)modename","pixels-per-inch",and"aspect-ratio"ifitisotherthanunity. ACKNOWLEDGEMENT Thisprogramusesthetype1hintingandrenderingsoftwareIBMhasdonatedtotheX-consor-tium. SEEALSO ‘METAFONT:TheProgram’,DonaldKnuth. ‘TheGFtoPKprocessor’,belongingtotheMETAFONTware. ‘AdobeFontMetricFiles’,Version3.0,PostScriptDeveloper SupportGroup. afm2tfm(1) pk2bm(1) TeX 2 PS2PK(1) PS2PK(1) VERSION 1.1(May1992) AUTHOR PietTutelaers rcpt@urc.tue.nl modifiedforps2pkm byPierreA.Mackay mackay@cs.washington.edu TeX 3